Grade 11Mechanics

A train starts from rest and moves with uniform acceleration ∆1 for some time and acquires velocity `v`. It then moves with constant velocity for some time and then decelerates at rate ∆2 and finally comes to rest at the next station. If `L` is distance between two stations then total time of travel is?

Yash
10 Years agoGrade 11

1 Answer

Vikas TU
10 Years ago
For ∆1 = a (let)
let it ttrravelled x distance for t1 time then,
x = 0.5at1^2
 
After,
y = v*t2
 
Then After,
z = vt3 – 0.5bt3^2 {let ∆2 = b}
also 0 = v^2 – 2bz
and given,
x + y + z = L
thus solve all the eqns and get the value of t1 and t2 and t3 and add them.